1.how to find average value of vector field if function of that field is given ?

Grade:12

1 Answers

Sandeep Kumar Rana
10 years ago
Consider field is given by some potential V=3x^2 and question is to find field vector at point x=3m? Ans then E=-dV/dx which becomes E=-6x=-6*3 since x=3 = -18V/m along -ve x direction